# Enhancing Multimodal In-Context Learning for Image Classification through Coreset Optimization

### v1

@@ Abstract.
In-context learning (ICL) enables Large Vision-Language Models (LVLMs) to adapt to new tasks without parameter updates, using a few demonstrations from a large support set. However, selecting informative demonstrations leads to high computational and memory costs. While some methods explore selecting a small and representative coreset in the text classification, evaluating all support set samples remains costly, and discarded samples lead to unnecessary information loss. These methods may also be less effective for image classification due to differences in feature spaces. Given these limitations, we propose Key-based Coreset Optimization (KeCO), a novel framework that leverages untapped data to construct a compact and informative coreset. We introduce visual features as keys within the coreset, which serve as the anchor for identifying samples to be updated through different selection strategies. By leveraging untapped samples from the support set, we update the keys of selected coreset samples, enabling the randomly initialized coreset to evolve into a more informative coreset under low computational cost. Through extensive experiments on coarse-grained and fine-grained image classification benchmarks, we demonstrate that KeCO effectively enhances ICL performance for image classification task, achieving an average improvement of more than 20%. Notably, we evaluate KeCO under a simulated online scenario, and the strong performance in this scenario highlights the practical value of our framework for resource-constrained real-world scenarios.
In-context Learning, Large Vision-Language Model, Coreset
ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
@@ 2.1. Large Vision-Language Model
Large Vision-Language Models (LVLMs) (Liang et al., 2024) are generally built on a vision encoder, a pre-trained Large Language Model (LLM), and an alignment module between the vision encoder and the LLM. The level of performance of these models has started to approach those of LLMs, especially after multimodal instruction tuning (Liu et al., 2023; Huang et al., 2023). Although there are numerous LVLMs available, it is important to note that not all of these models support in-context learning (ICL). For example, mPLUG-Owl (Ye et al., 2023), BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) lack the capabilities for ICL because they have not undergone dedicated few-shot pre-training and cannot handle the input distribution associated with ICL. In contrast, models like Flamingo (Alayrac et al., 2022) and IDEFICS (Laurençon et al., 2023) are specifically designed to support this task. In this work, we mainly focus on two open-source models with ICL capabilities (OpenFlamingo (Awadalla et al., 2023) and IDEFICS (Laurençon et al., 2023)) for our main experiments. Additional experiments are also performed on a high-performance commercial model, Qwen2-VL (Wang et al., 2024).
@@ 2.2. Multimodal In-context Learning
ICL enables LLMs to tackle novel tasks by leveraging a few demonstrations from a support set (Brown et al., 2020; Dong et al., 2022). Building on this success, ICL has been extended to LVLMs (Tsimpoukelli et al., 2021; Alayrac et al., 2022; Li et al., 2023e; Laurençon et al., 2024; Awadalla et al., 2023; Li et al., 2023f). To further enhance ICL in LVLMs, several works have focused on improving the construction strategies of in-context sequences, such as similarity-based selection (Liu et al., 2022; Pan, 2023; Li et al., 2023b; Min et al., 2022; Yang et al., 2023b) and diversity-based selection (Levy et al., 2023; Li et al., 2024). Despite its effectiveness, it faces challenges in resource-limited scenarios, requiring large storage and significant computational resources for computation (Yang et al., 2023a). Therefore, in the NLP domain, some works focus on selecting samples from an unlabeled pool for annotation to reduce annotation costs (Qian et al., 2024; Mavromatis et al., 2023), while others attempt to explore a representative coreset for ICL (Sener and Savarese, 2017; Li and Qiu, 2023). However, identifying such a coreset requires complex computation strategies and time-consuming evaluations of all samples in the support set. Additionally, a small coreset may not adequately represent the diverse semantics of the dataset in vision domain, and discarded data from the support set remains untapped. To overcome these limitations, we randomly sample a coreset and then integrate feature-level information from the untapped data to enhance the information contained within this coreset.
@@ 3.5. Key-based Sample Update
where $|S^{\prime}_{i,j}|$ is the number of samples associated with $t_{j}$, and $\alpha$ is a update rate that controls the size of the update step, which ranges from 0 to 1. This update strategy encourages each key to move toward the averaged features of its associated samples in a controlled manner.
@@ 3.6. Simulated Online Scenario
Online Target Sample Selection. To determine the target sample to update, we adopt the same selection strategies as in the general scenario: RS, SS, and DS, as described in Section 3.4.
@@ 3.7. Inference and Evaluation
Once we obtain the final coreset $\mathcal{C^{\prime}}$ by updating $\mathcal{C}$ via KeCO, we can evaluate a test image $\hat{\bm{I}}$ by retrieving the top-$k$ most relevant samples from $\mathcal{C}^{\prime}$, based on the similarity between $\phi(\hat{\bm{I}})$ and all stored keys $\bm{k}$. These retrieved samples constitute the in-context sequence $\mathcal{D}$, which is concatenated with the test image and fed into the frozen LVLM for prediction, as defined in Equation (2).
@@ 4. Experiments
$2$-shot $4$-shot $2$-shot $4$-shot $2$-shot $4$-shot $2$-shot $4$-shot $2$-shot $4$-shot $2$-shot $4$-shot
FS-IC $48.11$ $34.88$ $84.65$ $90.36$ 49.80 $50.22$ $74.42$ $82.25$ $65.26$ 62.92 90.52 94.98
FS-IS $61.04$ $58.61$ $85.62$ $93.09$ 61.42 $62.55$ $76.89$ $87.23$ $71.88$ 71.98 90.48 95.50
KeCO-SS $49.65$ $32.22$ $84.60$ $90.58$ 56.60 $51.54$ $75.59$ $82.27$ $70.66$ 63.60 90.78 95.88
KeCO-RS $72.90$ $72.66$ $85.45$ $94.34$ 72.20 $73.04$ $79.66$ $90.45$ $77.50$ 79.30 91.14 96.12
@@ 4.3. Implementation Details
We use top-1 accuracy as the metric to evaluate the performance due to its clarity and common usage. In general, each category in the coreset $C$ contains 5 or 10 samples (depending on the dataset), and the size $n$ of the support set $S$ is five times the size $m$ of $C$. Therefore, for CUB-200, $n=5000$ and $m=1000$; for Stanford Dogs, $n=6000$ and $m=1200$; and for ImageNet-100, $n=5000$ and $m=1000$. Since $S^{\prime}=S-C$, the sizes of $S^{\prime}$ are respectively 4,000, 4,800 and 4,000. Unless otherwise specified, the update rate $\alpha$ is set to 0.2, the epoch $e$ is set to 10, and the batch size $b$ is set to 1,000 (ablation study on $\alpha$, $e$ and $b$ are proveded in Appendix A.3).
@@
{tblr}
cells = c, column2 = fg=Gray, column3,5 = fg=black, column5 = fg=GuardsmanRed, hline1,4 = -0.08em, hline2 = -, Dataset & FS-IC FS-IS Online $\Delta$
@@ 5.2. Analyses of the Size of Additional Data and Coreset
Method OF-3B IDE-8B
1:2 1:4 1:6 1:2 1:4 1:6
FS-IC 49.80 49.80 49.80 74.42 74.42 74.42
FS-IS 58.67 61.42 63.42 75.68 76.89 77.45
KeCO-RD 69.49 72.20 73.58 78.32 79.66 79.13
KeCO-DS 69.99 73.75 74.58 79.02 79.72 79.92
Online Scenario
KeCO-RS 63.02 67.06 69.93 77.30 78.38 78.66
KeCO-DS 65.56 69.92 72.47 77.33 78.87 79.52
Method OF-3B IDE-8B
400 800 1000 400 800 1000
FS-IC 33.53 44.65 48.11 82.84 83.59 84.65
FS-IS 61.04 61.04 61.04 85.62 85.62 85.62
KeCO-RS 69.59 71.85 72.90 85.05 85.33 85.46
KeCO-DS 71.83 73.21 74.20 85.67 85.90 87.38
Online Scenario
KeCO-RS 65.75 67.86 65.90 84.04 84.51 85.69
KeCO-DS 69.47 71.92 70.61 85.38 86.69 85.91

### v2

@@ Abstract.
In-context learning (ICL) enables Large Vision-Language Models (LVLMs) to adapt to new tasks without parameter updates, using a few demonstrations from a large support set. However, selecting informative demonstrations leads to high computational and memory costs. While some methods explore selecting a small and representative coreset in the text classification, evaluating all support set samples remains costly, and discarded samples lead to unnecessary information loss. These methods may also be less effective for image classification due to differences in feature spaces. Given these limitations, we propose Key-based Coreset Optimization (KeCO), a novel framework that leverages untapped data to construct a compact and informative coreset. We introduce visual features as keys within the coreset, which serve as the anchor for identifying samples to be updated through different selection strategies. By leveraging untapped samples from the support set, we update the keys of selected coreset samples, enabling the randomly initialized coreset to evolve into a more informative coreset under low computational cost. Through extensive experiments on coarse-grained and fine-grained image classification benchmarks, we demonstrate that KeCO effectively enhances ICL performance for image classification task, achieving an average improvement of more than 20%. Notably, we evaluate KeCO under a simulated online scenario, and the strong performance in this scenario highlights the practical value of our framework for resource-constrained real-world scenarios. Our code is provided in: https://github.com/chenyil6/KeCO_Coreset_Optimization.
In-context Learning; Large Vision-Language Model; Coreset
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755764isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
@@ 2.1. Large Vision-Language Model
Large Vision-Language Models (LVLMs) (Liang et al., 2024) are generally built on a vision encoder, a pre-trained Large Language Model (LLM), and an alignment module between the vision encoder and the LLM. The level of performance of these models has started to approach those of LLMs, especially after multimodal instruction tuning (Liu et al., 2023; Huang et al., 2023; You et al., 2022; Das et al., 2024; Hu et al., 2024; You et al., 2021a, b; Peng et al., 2025). Although there are numerous LVLMs available, it is important to note that not all of these models support in-context learning (ICL). For example, mPLUG-Owl (Ye et al., 2023), BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) lack the capabilities for ICL because they have not undergone dedicated few-shot pre-training and cannot handle the input distribution associated with ICL. In contrast, models like Flamingo (Alayrac et al., 2022) and IDEFICS (Laurençon et al., 2023) are specifically designed to support this task. In this work, we mainly focus on two open-source models with ICL capabilities (OpenFlamingo (Awadalla et al., 2023) and IDEFICS (Laurençon et al., 2023)) for our main experiments. Additional experiments are also performed on a high-performance commercial model, Qwen2-VL (Wang et al., 2024).
@@ 2.2. Multimodal In-context Learning
ICL enables LLMs to tackle novel tasks by leveraging a few demonstrations from a support set (Brown et al., 2020; Dong et al., 2022). Building on this success, ICL has been extended to LVLMs (Tsimpoukelli et al., 2021; Alayrac et al., 2022; Li et al., 2023e; Laurençon et al., 2024; Awadalla et al., 2023; Li et al., 2023f; Feng et al., 2025). To further enhance ICL in LVLMs, several works have focused on improving the construction strategies of in-context sequences, such as similarity-based selection (Liu et al., 2022; Pan, 2023; Li et al., 2023b; Min et al., 2022; Yang et al., 2023b) and diversity-based selection (Levy et al., 2023; Li et al., 2024). Despite its effectiveness, it faces challenges in resource-limited scenarios, requiring large storage and significant computational resources for computation (Yang et al., 2023a; Jiang et al., 2025). Therefore, in the NLP domain, some works focus on selecting samples from an unlabeled pool for annotation to reduce annotation costs (Qian et al., 2024; Mavromatis et al., 2023), while others attempt to explore a representative coreset for ICL (Sener and Savarese, 2017; Li and Qiu, 2023). However, identifying such a coreset requires complex computation strategies and time-consuming evaluations of all samples in the support set. Additionally, a small coreset may not adequately represent the diverse semantics of the dataset in vision domain, and discarded data from the support set remains untapped. To overcome these limitations, we randomly sample a coreset and then integrate feature-level information from the untapped data to enhance the information contained within this coreset.
@@ 3.5. Key-based Sample Update
where $|S^{\prime}_{i,j}|$ is the number of samples associated with $t_{j}$, and $\alpha$ is a update rate that controls the size of the update step, which ranges from 0 to 1. This update strategy encourages each key to move toward the averaged features of its associated samples in a controlled manner, and the update process is repeated for $e$ epochs.
@@ 3.6. Simulated Online Scenario
Online Target Sample Selection.
To determine the target sample to update, we adopt the same selection strategies as in the general scenario: RS, SS, and DS, as described in Section 3.4.
@@ 3.7. Inference and Evaluation
[…]
@@ 4. Experiments
222-shot 444-shot 222-shot 444-shot 222-shot 444-shot 222-shot 444-shot 222-shot 444-shot 222-shot 444-shot
FS-IC 48.1148.1148.11 34.8834.8834.88 84.6584.6584.65 90.3690.3690.36 49.80 50.2250.2250.22 74.4274.4274.42 82.2582.2582.25 65.2665.2665.26 62.92 90.52 94.98
FS-IS 61.0461.0461.04 58.6158.6158.61 85.6285.6285.62 93.0993.0993.09 61.42 62.5562.5562.55 76.8976.8976.89 87.2387.2387.23 71.8871.8871.88 71.98 90.48 95.50
KeCO-SS 49.6549.6549.65 32.2232.2232.22 84.6084.6084.60 90.5890.5890.58 56.60 51.5451.5451.54 75.5975.5975.59 82.2782.2782.27 70.6670.6670.66 63.60 90.78 95.88
KeCO-RS 72.9072.9072.90 72.6672.6672.66 85.4585.4585.45 94.3494.3494.34 72.20 73.0473.0473.04 79.6679.6679.66 90.4590.4590.45 77.5077.5077.50 79.30 91.14 96.12
@@ 4.3. Implementation Details
We use top-1 accuracy as the metric to evaluate the performance due to its clarity and common usage. In general, each category in the coreset $C$ contains 5 or 10 samples (depending on the dataset), and the size $n$ of the support set $S$ is five times the size $m$ of $C$. Therefore, for CUB-200, $n=5000$ and $m=1000$; for Stanford Dogs, $n=6000$ and $m=1200$; and for ImageNet-100, $n=5000$ and $m=1000$. Since $S^{\prime}=S-C$, the sizes of $S^{\prime}$ are respectively 4,000, 4,800 and 4,000. Unless otherwise specified, the update rate $\alpha$ is set to 0.2, the epoch $e$ is set to 10, and the batch size $b$ is set to 1,000 (ablation study on $\alpha$, $e$ and $b$ are provided in Appendix A.3).
@@
Dataset FS-IC FS-IS Online $\Delta$
@@ 5.2. Analyses of the Size of Additional Data and Coreset
From Table 6 in Appendix A.4, it can be observed that as the amount of untapped data increases, there is an improvement in the FS-IS, KeCO-RS, and KeCO-DS settings. For instance, when the ratio of coreset size to untapped data size increases from 1:2 to 1:6, the performance of OF-3B under the FS-IS setting improves from 58.67% to 63.42%, and IDE-8B under the same setting improves from 75.68% to 77.45%. Similarly, under the KeCO framework, when the ratio changes from 1:2 to 1:6, OF-3B’s ICL performance under the KeCO-DS setting improves from 69.99% to 74.58%, and IDE-8B under the same setting improves from 79.02% to 79.92%. The analysis of the coreset size can be found in the Appendix A.4.
@@ 6. Conclusion
Acknowledgements.
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), and the Fundamental Research Funds for the Central Universities (2242025K30024). This research work is also supported by the Big Data Computing Center of Southeast University
@@ A.3. Ablation Study on Hyperparameters
A.4. Analyses of the Size of Additional Data and Coreset
Method OF-3B IDE-8B
1:2 1:4 1:6 1:2 1:4 1:6
FS-IC 49.80 49.80 49.80 74.42 74.42 74.42
FS-IS 58.67 61.42 63.42 75.68 76.89 77.45
KeCO-RD 69.49 72.20 73.58 78.32 79.66 79.13
KeCO-DS 69.99 73.75 74.58 79.02 79.72 79.92
Online Scenario
KeCO-RS 63.02 67.06 69.93 77.30 78.38 78.66
KeCO-DS 65.56 69.92 72.47 77.33 78.87 79.52
Method OF-3B IDE-8B
400 800 1000 400 800 1000
FS-IC 33.53 44.65 48.11 82.84 83.59 84.65
FS-IS 61.04 61.04 61.04 85.62 85.62 85.62
KeCO-RS 69.59 71.85 72.90 85.05 85.33 85.46
KeCO-DS 71.83 73.21 74.20 85.67 85.90 87.38
Online Scenario
KeCO-RS 65.75 67.86 65.90 84.04 84.51 85.69
KeCO-DS 69.47 71.92 70.61 85.38 86.69 85.91